An average weight of 3 containers A, B, C is 3.2kg. Container A is twice as heavy as Container B. Container B is 400g heavier that Container C. Find the weight of Container C.
Answer by ikleyn(52810)

.
An average weight of 3 containers A, B, C is 3.2 kg.
Container A is twice as heavy as Container B.
Container B is 400g heavier that Container C.
Find the weight of Container C.
~~~~~~~~~~~~~~~~~~

Let x = weight/(mass) of container C, in grams.

Then  container B weights (x+400) grams, while container A weights 2*(x+400) grams.


At this point, we write the total weight equation


        A    +   B     + C  = 3*3200       grams.    

    
    2(x+400) + (x+400) + x  = 9600 


Simplify and find x

    2x + 800 + x + 400 + x  = 9600

            4x              = 9600 - 800 - 400

            4x              =       8400

             x              =       8400/4 = 2100.


ANSWER.  Container C weights (has a mass of) 2100 grams.

Solved (with explanations and steps).